OpenStudy (anonymous):

Melanie’s monthly pay is $5553.60. Each month she uses 2/3 of her pay for household expenses and 1/10 for donations. She deposits the rest into her savings account. How much does Melanie deposit into her savings account each month?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

so you need to make 2/3 and 1/10 have the same denominator

OpenStudy (anonymous):

23/30?

OpenStudy (smored):

find\[\frac{ 2 }{ 3 }+\frac{ 1 }{ 10 }\]

OpenStudy (aripotta):

oh yes, 23/30 is what she paid for expenses and donations. so 1 - 23/30 or 30/30 - 23/30 should be what's left over

OpenStudy (anonymous):

so 7/1 or 7?

OpenStudy (aripotta):

30/30 - 23/30 = 7/30

OpenStudy (aripotta):

then you can multiply 7/30 and 5553.60 to get what was put into her savings

OpenStudy (anonymous):

.23*5553.60= 1277.328=1277.33

OpenStudy (aripotta):

it's not .23

OpenStudy (aripotta):

wait, sorry. you wanna do it in the calculator first tho to get a more exact answer. i got 1295.84

OpenStudy (aripotta):

don't round it tho, just keep it in your calculator and multiply from there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I did it the way you told me and got the same thing

OpenStudy (aripotta):

so type in: 7 / 30 = x 5553.60 =

OpenStudy (anonymous):

Thank you!

OpenStudy (aripotta):

like don't click clear after you do 7/30

OpenStudy (aripotta):

no problemo :)

OpenStudy (anonymous):

1295.84 is the correct answer!